OpenStudy (anonymous):

two integers differ by 12 and the sum of their squares is 74. find the integers two integers differ by 12 and the sum of their squares is 74. find the integers @Mathematics

OpenStudy (ajprincess):

Let the integers be x and y. x-y=12 x=y+12 x^2+y^2=74 (y+12)^2+y^2=74 y^2+24y+144+y^2=74 2y^2+24y+144-74=0 2y^2+24y+70=0 y^2+12y+35=0 y^2+7y+5y+35=0 y(y+7)+5(y+7)=0 (y+7)(y+5)=0 y+7=0 or y+5=0 y=-7 or y=-5 plug in the values of y in x-y=12 to find x. There will be two values